Aaronsssssss:

Nicholas uses a credit card with a 19.1% APR, compounded monthly, to pay for a cruise totaling $2,513.69. He can pay $910 per month on the card. What will the total cost of this purchase be?

izuwu1:

is the credit card his?

Aaronsssssss:

Yes

jimthompson5910:

We have 19.1% APR, which is the annual rate What is the monthly rate?

Aaronsssssss:

$2.993.80 $2,513.69 $3,285.64 $2,591.95 these are the choices it gives

jimthompson5910:

hint: monthly rate = (APR)/12

Aaronsssssss:

So 910=(.191)/12

Aaronsssssss:

?

jimthompson5910:

what does your calculator say when you type in 0.191/12

Aaronsssssss:

0.015

jimthompson5910:

I'm getting 0.01591666666667 Let's round that to 0.0159167 just to keep things relatively simple

jimthompson5910:

now multiply that with 2,513.69 and tell me what you get

Aaronsssssss:

40.0095658

jimthompson5910:

I'm getting the same. Let's round that to 40.01 since money is usually rounded to the nearest cent (hundredth)

Aaronsssssss:

Okay

jimthompson5910:

The 40.01 is the amount of interest for the month The balance 2513.69 jumps up to 2553.70 after we add on the 40.01 In other words, 2513.69+40.01 = 2553.70

jimthompson5910:

So we started with a balance of 2513.69 and now the balance is 2553.70

Aaronsssssss:

Correct

jimthompson5910:

Then the payment of $910 will lower the balance to 2553.70-910 = 1643.70 With me so far?

Aaronsssssss:

Yes

jimthompson5910:

ok that was all for month #1

jimthompson5910:

we will repeat this process for month #2

jimthompson5910:

start with a balance of 1643.70 and find the monthly interest you'll pay So you'll need to multiply 1643.70 with 0.0159167

Aaronsssssss:

26.16227979

jimthompson5910:

that rounds to 26.16

jimthompson5910:

Add that onto the balance and you get what?

Aaronsssssss:

1,669.86?

jimthompson5910:

correct

jimthompson5910:

now subtract off the monthly payment of $910

Aaronsssssss:

759.86

jimthompson5910:

correct we'll need to do one more round of this

Aaronsssssss:

Ok

jimthompson5910:

if the balance is 759.86, what is the monthly interest we pay for month #3?

Aaronsssssss:

How do I get that

jimthompson5910:

multiply 759.86 with the value 0.0159167

jimthompson5910:

since 0.0159167 was the monthly rate (in decimal form)

jimthompson5910:

monthly interest = (current balance)*(decimal form of monthly rate)

Aaronsssssss:

Multiply by 759?

jimthompson5910:

759.86 * 0.0159167 = ??

Aaronsssssss:

12.09

jimthompson5910:

yes, so the balance goes from 759.86 to 759.86+12.09 = 771.95 Agreed?

Aaronsssssss:

Yes

jimthompson5910:

At this point, we would subtract off the payment of $910 However, this payment is larger than the current balance So we can just pay off the entire balance and bring it to 0 This means the final payment is $771.95 instead of $910

jimthompson5910:

Throughout this process, you should be keeping track of all the values in a chart like you see below

Aaronsssssss:

Oh okay

jimthompson5910:

The table is handy to see how much you pay on a given month, and what your balance is, etc

jimthompson5910:

what do you get when you add up the stuff in the "payment" column?

Aaronsssssss:

So add everything?

jimthompson5910:

just in the payment column

jimthompson5910:

doing so will lead to the total cost, since it's the sum of all the payments

Aaronsssssss:

2,591.95

jimthompson5910:

correct

jimthompson5910:

so the answer is choice D

Aaronsssssss:

Thank you

jimthompson5910:

If you're curious about the total interest paid, then you subtract the total cost and cost of the product 2591.95 - 2513.69 = 78.26 The total interest is $78.26

jimthompson5910:

no problem